OpenStudy (rayford):

What is the value for the coefficient determinant (D) in the following system? 2x + y = 2 3x + 4y = –7 [a]1 [b]5 [c]11 [d]15

OpenStudy (anonymous):

For the coefficient determinant, make the coefficients of x one column and the coefficients of y the second.

OpenStudy (anonymous):

So it would be like this: First Row (2, 1) Second Row(3, 4)

OpenStudy (rayford):

i know, i got to that answer but i dont know which one it is?

OpenStudy (anonymous):

To evaluate the determinant, use this formula for determinant with Row 1: (a, b) Row 2 (c, d) Det = ad - bc So for this problem = 2*4 - 3*1 = 5

OpenStudy (anonymous):

So its B..

OpenStudy (rayford):

thank you so much :]

OpenStudy (anonymous):

your welcome..

OpenStudy (anonymous):

@Trexy ok its five but what about right side?